OpenStudy (anonymous):

12x+9=-4x^2 The discriminant is: a. 288 b. 108 c. -132 d. 0

OpenStudy (anonymous):

b. 108

OpenStudy (anonymous):

Thanks! How'd you get the answer?

OpenStudy (anonymous):

no it will be 0

OpenStudy (anonymous):

Oh o.k. Can you please explain? I have no clue how to solve :/

OpenStudy (anonymous):

Kreshnik were you typing? Nothing showed up :/

OpenStudy (anonymous):

\[\Large 12x+9=-4x^2\] \[\Large 4x^2+12x+9=0\] now here you got the quadratic equation... use the formula.. \[\Large x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] substitute a=4 b=12 c=9 and you'll get... \[\Large \Delta =b^2-4ac\] \[\Large \Delta =144-4\cdot 4 \cdot 9\] \[\Large \Delta =144-144\] \[\Large \ \Delta =0\] so when Delta =0 we have only one solution... and going on with the formula the solution will be \[\Large x=-\frac32 \quad \quad or \quad \quad x=-1.5\] NONE OF THE MULTIPLE CHOICES ... @Rohangrr Stop giving blindly answers !

OpenStudy (anonymous):

Whoa thanks Kreshnik! Awesome explanation

OpenStudy (anonymous):

@mathsux4real Glad to help. :) ... You're welcome ;)